OpenStudy (anonymous):

complete the square. x^2 - 6x= -8

OpenStudy (anonymous):

x^2 - 6x= -8 x^2 - 2(3)x+8+1-1=0 x^2 - 2(3)x+9=1 (x+3)^2=1

OpenStudy (anonymous):

here you can add (b/2)^2 to both sides

OpenStudy (anonymous):

(x-3)^2=1

OpenStudy (anonymous):

@matricked @minimoo ,(x-3)^2=1 ,x^2 - 2(3)x+9-1=0,then x^2-6x+8=0,so x^2-6=-8

OpenStudy (anonymous):

guys ,the solutions of this equation are 4 and 2

OpenStudy (anonymous):

sorry x -3

OpenStudy (anonymous):